### v1

# references.bib

Abstract
Autonomous Vehicles (AVs) collect and pseudo-label terabytes of multi-modal data localized to HD maps during normal fleet testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs remains a significant challenge. Traditional scenario mining techniques are error-prone and prohibitively time-consuming, often relying on hand-crafted structured queries. In this work, we revisit spatio-temporal scenario mining through the lens of recent vision-language models (VLMs) to detect whether a described scenario occurs in a driving log and, if so, precisely localize it in both time and space. To address this problem, we introduce RefAV, a large-scale dataset of $10,000$ diverse natural language queries that describe complex multi-agent interactions relevant to motion planning derived from $1000$ driving logs in the Argoverse 2 Sensor dataset. We evaluate several referential multi-object trackers and present an empirical analysis of our baselines. Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges. Our code and dataset are available on Github and Argoverse.
1 Introduction
Autonomous Vehicle (AV) deployment on public roads has increased significantly in recent years, with Waymo completing more than a million rides per week as of February 2025 [Korosec2025]. Despite the maturity of such autonomous ride-hailing services, AVs still get into accidents [thadani2023cruise] and often require manual intervention [korosec2025waymo, mihalascu2023cruise]. Although data-driven simulators are an integral component in establishing a safety case [karnchanachari2024towards, gulino2023waymax, vasudevan2024planning, yang2023unisim], validating end-to-end autonomy with real-world operational data remains a critical part of testing. However, identifying interesting and safety-critical scenarios from uncurated driving logs is akin to finding a "needle in a haystack" due to the scale of data collected during normal fleet operations. In this paper, we revisit the task of spatio-temporal scenario mining with recent vision-language models (VLMs) to identify interesting multi-agent interactions using natural language (cf. Fig. 1).
Status Quo. Although language-based 3D scene understanding has been extensively studied in the context of referential multi-object tracking (RMOT) [wu2023referkitti, wu2023nuprompt, li2025nugrounding], multi-modal visual question answering (VQA) [qian2024nuscenesqa, Inoue_2024_WACV, tian2025nuscenes], and VLM-based motion planning [sima2024drivelm, mao2023gpt, shao2024lmdrive, wang2024omnidrive], we argue that scenario mining presents unique challenges. RMOT extends referential grounding by associating referred objects over time. However, unlike RMOT, scenario mining does not guarantee that referred objects exist in a given driving log. Next, multi-modal VQA extends VQA with additional visual modalities like LiDAR. Although spatio-temporal scenario mining also leverages multi-modal input (e.g. RGB, LiDAR and HD maps), we require that methods output 3D tracks instead of text-based answers. Lastly, VLM-based motion planners directly estimate future ego-vehicle waypoints based on high-level language instructions and past sensor measurements. In contrast, scenario mining methods can reason over the full driving log to identify interactions between non-ego vehicles. Concretely, spatio-temporal scenario mining requires identifying if a described scenario occurs in a driving log from raw sensor measurements (e.g. RGB and LiDAR), and if so, precisely localizing all referred objects in both time and space with 3D tracks. To support this task, we propose RefAV, a large-scale dataset of $10,000$ diverse natural language queries designed to evaluate a model’s ability to find a visual “needle in a haystack”.
RefAV Scenario Mining Dataset. We repurpose the Argoverse 2 (AV2) Sensor dataset [wilson2023argoverse], which contains $1,000$ driving logs with synchronized LiDAR, 360 ring cameras, HD maps, and 3D track annotations for $30$ categories. We curate a set of $10,000$ natural language prompts that describe interesting (and often rare) scenarios (cf. Fig. 2 and Fig. 4) using a combination of manual annotations and procedural generation using large language models (LLMs). Notably, AV2 annotates ground truth tracks at 10 Hz, which allows for fine-grained motion understanding. In contrast, prior referential multi-object tracking benchmarks (the established task most similar to spatio-temporal scenario mining) use nuScenes [caesar2020nuscenes], which annotates ground truth tracks at 2 Hz (cf. Table 1). Argoverse’s higher temporal resolution uniquely allows RefAV to evaluate dynamic multi-agent interactions (unlike prior work that primarily evaluates referential expressions based on static attributes like vehicle color or relative heading). Interestingly, we find that simply repurposing VLMs for scenario mining yields poor performance.
Figure 1: Scenario Mining Problem Setup. Given a natural language prompt such as vehicle making left turn through ego-vehicle’s path while raining, our problem setup requires models to determine whether the described scenario occurs within a 20-second driving log, and if so, precisely localize the referred object in 3D space and time from raw sensor data (LiDAR, 360 ring cameras, and HD maps). Based on the example above, a VLM should localize the start and end timestamps of the red Mini Cooper executing a “Pittsburgh left” through the ego-vehicle’s path with a 3D track. Notably, the “Pittsburgh left” is a regional driving practice where a driver quickly makes a left turn before oncoming traffic proceeds. Although common in Pittsburgh, this maneuver is technically illegal. Therefore, we argue that scenario mining is critical for validating end-to-end autonomy in order to build a comprehensive safety case. Note that referred objects are shown in blue, related objects in green, and other objects in red.
Referential Tracking by Program Synthesis. Although prior referential trackers and VLMs achieve reasonable performance with simple referential prompts (e.g. find all cars), we find that such methods struggle with compositional reasoning and motion understanding (e.g. find all cars accelerating while changing lanes). To address this problem, we propose Referential Tracking by Program Synthesis (RefProg), a modular approach that combines off-the-shelf 3D tracks with LLMs. Inspired by recent work in program synthesis [surismenon2023vipergpt, gupta2023visual, marsili2025visual], our method uses an LLM to break down complex referential expressions into simpler compositional actions. Specifically, we define an API to describe hand-crafted atomic functions (e.g., turning, accelerating, changing lanes) and use an LLM to synthesize a program that composes these atomic actions corresponding to the natural language prompt. We then execute the generated program to filter off-the-shelf tracks to identify the subset of tracks that best align with the described scenario (cf. Fig. 3).
Contributions. We present three major contributions. First, we introduce RefAV, a large-scale dataset designed to evaluate VLMs on 3D scene understanding and spatio-temporal localization. Our extensive experiments highlight the limitations of current methods, and demonstrate the effectiveness of our proposed program synthesis-based approach. Lastly, we host a challenge at CVPR 2025 in conjunction with the Workshop on Autonomous Driving to encourage broad community involvement in addressing this challenging problem.
2 Related Works
Referential Grounding and Tracking are long-standing challenges in vision-language understanding. Early datasets like RefCOCO, RefCOCO+, and RefCOCOg [refcoco, refcoco+, refcocog] helped popularize the task, which was later adapted for autonomous driving [deruyttere2019talk2car, Vasudevan_2018_CVPR]. Although prior work focused on single-frame visual grounding, ReferKITTI [wu2023referkitti] formalizes the problem of 2D referential multi-object tracking (RMOT), which extends referential object detection by associating referred objects over time. ReferKITTI-v2 [zhang2024bootstrapping] expands ReferKITTI’s manual annotations using large language models (LLMs) to improve prompt diversity. Further, LaMOT [li2024lamot] consolidates 1,660 sequences from four tracking datasets to create a large-scale unified RMOT benchmark.
Referential tracking methods can be broadly classified into two-stage (e.g. separately track and filter predictions based on natural language cues) and end-to-end approaches. iKUN [du2024ikun] proposes a plug-and-play knowledge unification module to facilitate language grounding for off-the-shelf trackers. Similarly, ReferGPT [chamiti2025refergpt] presents a zero-shot method for filtering off-the-shelf 3D tracks using CLIP [clip] scores and fuzzy matching. In contrast, JointNLT [zhou2023joint] proposes a unified visual grounding and tracking framework, and MMTrack [mmtrack] reformulates tracking as token generation. OVLM [patzold2025leveraging] introduces a memory-aware model for RMOT, while ReferFormer [wu2022language] leverages language-conditioned object queries. More recently, nuPrompt [wu2023nuprompt] extends 2D referential tracking to 3D with RGB input. Different from nuPrompt, RefAV addresses 3D multi-modal referential tracking and dynamic scene understanding.
Visual Question Answering with Vision-Language Models has improved significantly in recent years due to large-scale multi-modal pre-training. Vision-language models (VLMs) like LLaVA [liu2023llava, liu2023improvedllava], BLIP-2 [blip2], and Qwen2.5-VL [bai2025qwen2.5vl] show strong generalization across diverse domains [balanced_vqa_v2, tian2025nuscenes, qian2024nuscenesqa]. However, state-of-the-art models still struggle with spatial reasoning and grounding [madan2024revisiting, robicheaux2025roboflow100]. To address this issue, SpatialRGPT [cheng2024spatialrgpt] and SpatialVLM [chen2024spatialvlm] generate large-scale monocular depth pseudo-labels to improve 3D reasoning. Despite the effectiveness of such methods, zero-shot prompting yields poor compositional reasoning performance. Instead, VisProg [gupta2023visual] and ViperGPT [surismenon2023vipergpt] use LLMs generate executable code and use tools like OwlViT [minderer2022simple], CLIP [clip], MiDaS [ranftl2020towards], and GLIP [li2022grounded] for spatial understanding. However, existing program synthesis approaches primarily focus on single-frame reasoning. In contrast, we are interested in understanding dynamic multi-agent interactions from video sequences. We take inspiration from existing methods and apply similar program synthesis-based approaches for spatio-temporal scenario mining.
Vision-Language Models for 3D Understanding have been extensively explored in the context of representation learning, open-vocabulary 3D perception, and end-to-end driving. SLidR [sautier2022image] distills 2D features into 3D point clouds for cross-modal learning. SEAL [liu2023segment] incorporates SAM [kirillov2023segment] to produce class-agnostic 3D segments, while SA3D [cen2024segment] leverages SAM and NeRFs for object segmentation. Similarly, recent work like Anything-3D [shen2023anything] and 3D-Box-Segment-Anything integrate VLMs and 3D detectors (e.g., VoxelNeXt [chen2023voxelnext]) for interactive reconstruction and labeling. VLMs have been used extensively in open-vocabulary perception for autonomous driving. UP-VL [najibi2023unsupervised] distills CLIP features into LiDAR data to generate amodal cuboids. Recent work uses 2D VLMs to generate 3D pseudo-labels for open-vocabulary perception, enabling zero-shot LiDAR panoptic segmentation [sal2024eccv, takmaz2025cal] and 3D object detection [khurana2024shelf, peri2022towards, ma2023long]. While traditional grounding methods [wu2023nuprompt] struggle with complex instructions, multi-modal LLMs [mao2023gpt, shao2024lmdrive, cho2024language] demonstrate impressive visual understanding. However, such methods [bai20243d, wang2024omnidrive, ding2024holistic] typically produce scene-level analysis for end-to-end driving rather than precise instance localization. In contrast, our framework combines language grounding with precise geometric localization from offline 3D perception methods for more accurate vision-language reasoning.
3 RefAV: Scenario Mining with Natural Language Descriptions
In this section, we present our approach for curating $10,000$ natural language prompts (Sec 3.1) and describe four zero-shot scenario mining baselines (Sec 3.2).
Figure 2: RefAV Dataset Creation. First, we define a set of 28 atomic functions that identify the state of an object track, its relationship with other objects (e.g. a scene graph), and a set of boolean logical operators to support function composition. Next, we prompt an LLM to permute these atomic functions and generate a program and corresponding natural language prompt. We prompt a separate LLM with the generated program and natural language prompt to verify the code. We additionally manually verify the code for correctness. Lastly, we execute the generated code on ground-truth tracks and visualize the referred object track to manually verify that the program output matches the natural language prompt. We sample valid programs to maximize scenario diversity in our dataset. Note that yellow boxes highlight automated processes, while blue boxes indicate steps requiring manual intervention.
3.1 Creating RefAV
Unlike prior benchmarks that primarily evaluate referential expressions based on static attributes like vehicle color or relative heading (e.g. find the red car to the left), we focus on mining interesting and safety-critical scenarios relevant for motion planning (cf. Fig. 1). We take inspiration from recent planning benchmarks like nuPlan [karnchanachari2024towards] to identify such planning-centric scenarios. In particular, nuPlan introduces a set of 80 scenarios considered relevant for safe motion planning. We use this as a template to identify similar scenarios within AV2.
Why Use the Argoverse 2 (AV2) Sensor Dataset? The AV2 sensor dataset contains $1000$ 15 – 20 second logs with synchronized sensor measurements from seven seven ring cameras and two LiDAR sensors. Moreover, the dataset includes HD maps with lane markings, crosswalk polygons, and various lane types (e.g. vehicle, bus, and bike lanes). Notably, AV2 annotates 30 object categories with track-level annotations at 10 Hz. We choose to build RefAV on top of AV2 because prior datasets like nuScenes [caesar2020nuscenes] only annotate ground-truth tracks at a lower temporal resolution (e.g. 2Hz vs. 10Hz), making it more difficult to extract fine-grained motion. We ablate this in Appendix D. In addition, KITTI [geiger2012we] and Waymo Open Dataset [sun2020scalability] only label a limited number of categories (e.g. car, pedestrian, bicycle) and do not have HD maps, making it difficult to evaluate diverse multi-agent interactions. Although AV2 annotates objects up to 150m away from the ego vehicle, we clip all object tracks at 50m. We find that current 3D perception models struggle with long-range detection and tracking [peri2023empirical], suggesting that the community is not ready to address spatio-temporal scenario mining at range.
LLM-based Procedural Scenario Generation. Our key insight is that many complex scenarios (e.g. find all cars accelerating while changing lanes) can be broken down into simpler atomic actions (e.g. find cars, accelerating, and changing lanes), and a large set of such atomic actions can be composed to generate new diverse scenarios. To this end, we first define 28 atomic functions based on nuPlan’s list of planning scenarios (cf. Figure 2 left). We include a full API listing in Appendix F. To generate a new scenario definition, we provide an LLM (e.g. Claude 3.5 Sonnet) with the full API listing, along with in-context examples of real compositions. We prompt the LLM to generate permutations of the atomic functions and describe the generated code with a natural language description (cf. Figure 2 middle). We execute the generated code to filter ground-truth tracks from all $1000$ logs to identify true positive matches. We aggregate all true positive log-prompt matches (>50K) and sample 8,000 true positive log-prompt pairs that maximize scenario diversity. Lastly, we randomly sample 2,000 true negative log-prompt pairs.
Verifying Procedurally Generated Scenarios. Although LLM-based procedural generation allows us to generate diverse scenarios at scale, this process is not perfect and requires extensive manual validation (cf. Figure 2 right). First, we verify that the generated code aligns with the generated natural language description. We then manually review video clips of true positive log-prompt matches to ensure that the generated natural language prompt accurately describes the localized tracks. Notably, we find two common error modes. First, the LLM often reverses the relationship between objects in the code and prompt. For example, given a prompt to find bicycles in front of the car, the generated code might actually describe the scenario car in front of the bicycles. Second, we find that LLMs often generates underspecified code. For example, given the prompt bicycles traveling against traffic, the LLM correctly generates code that finds bicycles riding in the opposite direction of the ego vehicle, but it does not precisely specify that they should be on the same side of the road as the ego vehicle.
Manual Scenario Annotation. Despite the versatility of our procedural generation approach, many interesting interactions cannot be easily identified by composing atomic functions. Therefore, we manually inspect all videos in the validation and test sets to identify interesting driving behaviors (cf. Fig 4). Further, we manually annotate weather (e.g. clear, cloudy, rain, snow, and fog) and lighting conditions (e.g. daylight, dusk/dawn, and night) in the validation and test sets as these attributes are relevant for establishing a safety case [lee2024typical] in all driving conditions. See Appendix B for more details about our annotation workflow.
Dataset Statistics. As shown in Table 1, our dataset uniquely addresses the task of spatio-temporal scenario mining, whereas other datasets focus on object detection, referring multi-object tracking (RMOT), or visual question answering (VQA). Further, RefAV is uniquely built on Argoverse 2 [wilson2023argoverse], whereas most existing datasets rely on nuScenes [caesar2020nuscenes] or KITTI [geiger2012we]. Next, RefAV provides track-level annotations at a higher frequency of 10Hz, compared to the typical 2Hz in other datasets, allowing for more fine-grained temporal analysis. While datasets like nuGrounding [li2025nugrounding] and nuScenes-QA [qian2024nuscenesqa] include large numbers of expressions, RefAV emphasizes a diversity of annotation types, including referring expressions that capture dynamic multi-agent interactions, weather, and lighting conditions. Lastly, RefAV is one of only two datasets to support negative prompts.
Table 1: Comparison to Other Benchmarks. Language-based 3D scene understanding has been extensively studied in the context of referential multi-object tracking (RMOT) and multi-modal visual question answering (VQA). Different from prior work, we address the problem of spatio-temporal scenario mining. Specifically, RefAV is based on Argoverse 2, which provides 3D track-level annotations for 30 categories at 10 Hz. Although RefAV does not include as many referential expressions as prior work (e.g. OmniDrive [wang2024omnidrive] and nuGrounding [li2025nugrounding]), our referential annotations focus on capturing diverse multi-agent interactions. Lastly, RefAV includes negative prompts, which allows us to more accurately measure scenario mining performance.
Dataset Base Data Task View Anno. Freq. # Expressions # Frames Neg. Prompts Human Anno.
Talk2Car [deruyttere2019talk2car] nuScenes Detection Front Per Video 11959 9217 Referring Expression
Refer-KITTI [wu2023referkitti] KITTI RMOT Front 2Hz 818 6650 Referring Expression
nuGrounding [li2025nugrounding] nuScenes RMOT 360 2Hz 2.2M 34149 Object Attribute
nuPrompt [wu2023nuprompt] nuScenes RMOT 360 2Hz 40776 34149 Object Color
nuScenes-QA [qian2024nuscenesqa] nuScenes VQA 360 2Hz 459941 34149 None
DriveLM [sima2024drivelm] nuScenes VQA Front 2Hz 443K 4871 QA Pairs
OmniDrive [wang2024omnidrive] nuScenes VQA 360 2Hz 200K 34149 None
RefAV (Ours) Argoverse 2 Scenario Mining 360 10Hz 10K 15500
Referring Expression,
Weather, Lighting
3.2 Scenario Mining Baselines
Figure 3: Method Overview. RefProg is a dual-path method that independently generates 3D perception outputs and Python-based programs for referential grounding. Given raw LiDAR and RGB inputs, RefProg runs a offline 3D perception model to generate high quality 3D tracks. In parallel, it prompts an LLM to generate code to identify the referred track. Finally, the generated code is executed to filter the output of the offline 3D perception model to produce a final set of referred objects (shown in green), related objects (shown in blue), and other objects (shown in red).
We present four referential tracking baselines for scenario mining, including filtering by referred class, referential track inflation, referential track classification, and referential tracking by program synthesis. We repurpose 3D trackers from winning teams in the Argoverse 2 End-to-End Forecasting Challenge.
Filtering by Referred Class (RefBlind). We propose a simple “blind” baseline that does not explicitly reason about multi-agent interactions. For a given natural language prompt, we use an LLM to parse the referred object class and only keep 3D tracks from this class. For example, for the natural language query find all cars turning left, we remove all predicted tracks except cars. Interestingly, RefBlind is a surprisingly strong baseline, even compared to more complex approaches, described next.
Referential Track Inflation (RefInflate). We take inspiration from CM3D [khurana2024shelf] to create a zero-shot referential tracker using off-the-shelf 2D foundation models. First, we prompt an 2D detector with referential grounding capabilities (e.g., GroundingDINO [liu2024grounding]) with a natural language query (e.g., car turning left) to generate 2D box proposals. Next, we prompt SAM [kirillov2023segment] with the predicted 2D bounding boxes to generate high-quality instance segmentation masks. We then generate an oriented 3D cuboid using the set of LiDAR points that project to the 2D instance mask. Lastly, we associate these referred 3D detections over time using a 3D Kalman filter [weng20203d].
Referential Track Classification (RefClassify). We take inspiration from ReferGPT [chamiti2025refergpt] to classify the output of off-the-shelf 3D trackers using CLIP [clip] into referred objects and other objects. For each track, we project the predicted 3D bounding box onto the 2D image plane and extract CLIP image features at each timestep. Next, we compute the cosine similarity between the per-timestep CLIP image embeddings and the CLIP text embedding of the referential prompt. We consider a track to be a referreed object if the similarity score between the image embeddings and text embedding is above a fixed threshold for more than 50% of the track length.
Referential Tracking by Program Synthesis (RefProg). RefProg takes a similar approach as our LLM-based procedural scenario generator. Specifically, we prompt an LLM to generate code using an API listing of atomic actions for a given referential query. We execute the generated code on off-the-shelf 3D tracks to identify referred objects, related objects and other objects (cf. Fig 3). Despite the similarity of RefProg and our LLM-based procedural scenario generator, we highlight three key differences. First, RefProg generates code based on the referential query, while our LLM-based procedural scenario generator synthesizes the referential query based on the generated code. We posit that the role of the scenario generator is considerably easier since there are many valid natural language prompts to describe a generated program, but significantly fewer valid programs corresponding to each natural language prompt. Second, RefProg filters predicted 3D tracks, while our scenario generator uses ground-truth 3D tracks. Lastly, code from our scenario generator is verified for correctness, while code from RefProg is used without modification.
4 Experiments
In this section, we briefly describe our evaluation metrics (Sec. 4.1), provide an empirical analysis of our baselines (Sec. 4.2), and ablate the impact of different LLMs on RefProg’s spatio-temporal scenario mining accuracy (Sec. 4.3).
4.1 Metrics
We evaluate scenario mining methods using HOTA-Temporal and HOTA-Track (variants of HOTA [luiten2021hota]) to measure referential tracking performance and balanced accuracy to measure information retrieval accuracy. HOTA is a unified metric that explicitly balances detection accuracy, association, and localization, making it better aligned with human visual evaluations of tracking performance. HOTA-Temporal extends the standard HOTA metric by only considering the referred timestamps within a track as true positives. HOTA-Track is similar to HOTA-Temporal, but it does not penalize methods for incorrectly predicting the start and end of a referred action. For example, given the prompt car turning right, HOTA-Temporal only considers the timestamps where the car is turning right in the full referred track as true positives, whereas HOTA-Track considers all timestamps in the referred track as true positives. We also include HOTA results to contextualize standard tracking performance with the referential tracking performance of all benchmarked trackers. Additionally, we use timestamp balanced accuracy as a timestamp-level classification metric and log balanced accuracy as a log-level classification metric to evaluate how well methods can identify which timestamps and logs contain objects that match the referential prompt, respectively.
For each prompt, we categorize all objects into three groups: referred objects, which are the primary objects specified by the prompt; related objects, which are objects that interact with the referred object; and other objects, which are neither referred to nor relevant to the prompt. We compute both HOTA-Temporal and HOTA-Track metrics exclusively for the referred object class in the main paper. We report the performance on related objects and other objects in Appendix C.
Table 2: Experimental Results. We evaluate several zero-shot referential tracking baselines. We find that RefProg significantly outperforms all other zero-shot baselines. We posit that this can be attributed to both RefGen and RefClass not effectively leveraging temporal context.
Method HOTA $\uparrow$ HOTA-Temporal $\uparrow$ HOTA-Track $\uparrow$ Log Bal. Acc. $\uparrow$ Timestamp Bal. Acc. $\uparrow$
Ground Truth 100.0 13.3 20.5 50.0 50.0
Filtering by Referred Class (RefBlind)
Ground Truth 100.0 21.4 33.9 52.9 57.2
LE3DE2E [chenle3de2e] 74.4 19.2 30.0 53.4 55.0
Referential Track Inflation (RefInflate)
GroundingDINO [liu2024grounding] - 4.0 6.0 49.7 49.3
Referential Track Classification (RefClassify)
Ground Truth 100.0 24.6 28.3 54.7 57.0
Le3DE2E [chenle3de2e] 74.4 17.2 24.4 51.1 51.1
ReVoxelDet [lee2024revoxeldet] 63.1 17.1 21.2 52.3 54.4
TransFusion [bai2022transfusion] 63.6 16.0 18.2 52.8 55.3
Referential Tracking by Program Synthesis (RefProg)
Ground Truth 100.0 64.8 68.7 81.1 80.7
Le3DE2E [chenle3de2e] 74.4 50.1 51.1 71.8 74.6
ReVoxelDet [lee2024revoxeldet] 63.1 43.9 45.1 70.7 69.8
TransFusion [bai2022transfusion] 63.6 48.1 46.6 70.7 68.7
Valeo4Cast [xu2024valeo4cast] 69.2 41.7 44.1 62.7 69.1
4.2 Empirical Analysis of Results
Table 2 presents several simple baselines for addressing spatio-temporal scenario mining. We evaluate ground-truth oracles for each zero-shot referential tracker to understand how tracking accuracy and language grounding independently influence final performance. First, evaluating all ground truth tracks without considering language prompts yields a HOTA-Temporal of 13.3%. Further, evaluating RefBlind with ground truth tracks improves referential tracking accuracy by 8%. However, RefBlind with LE3DE2E tracks results in a slight performance drop. Next, we find that RefInflate performs the worst out of all four baselines, as inflating 2D detections to 3D is challenging. In contrast, RefClassify performs considerably better. Interestingly, the RefClassify oracle outperforms the RefBlind oracle, suggesting that CLIP-based filtering can capture richer semantics than class names alone. However, RefClassify underperforms RefBlind when using predicted tracks (e.g., LE3DE2E scores 17.2 vs. 19.2 HOTA-Temporal). Finally, our proposed RefProg baseline significantly outperforms all other zero-shot methods. Notably, RefProg with LE3DE2E tracks achieves a 32.8% improvement in HOTA-Temporal over RefClassify with Le3DE2E tracks.
Analysis of Failure Cases. We find that all three baselines have unique failure modes. First, we note that RefInflate often predicts bounding boxes even in cases where no referred object exists. We attribute this to GroundingDINO’s [liu2024grounding] referential grounding pre-training, which biases the model towards always predicting a bounding box since each training image always has at least one true positive grouth-truth box. Next, we find that RefClassify performs poorly because the CLIP image embedding for each track lacks the context required to understand multi-agent interactions. Specifically, the 2D projected image only includes the tracked object, and lacks temporal context. Lastly, we find that RefProg fails in cases where the API listing cannot generate a valid program for a given prompt (e.g. prompts involving weather and lighting conditions). Although one can always add new atomic actions, this strategy is not scalable.
Table 3: Impact of Different LLMs on Code Synthesis Quality. We evaluate the impact of using different LLMs for program synthesis in the RefProg pipeline. Interestingly, we find that Claude 3.5 Sonnet has the lowest program failure rate (e.g. 99.5% of Claude 3.5 Sonnet’s generated programs were valid, compared to 81.9% for Qwen 2.5B Instruct), while Claude 3.7 Sonnet achieves the highest HOTA-Temporal.
LLM Failure Rate $\downarrow$ HOTA-Temporal $\uparrow$ HOTA-Track $\uparrow$ Log Bal. Acc. $\uparrow$ Timestamp Bal. Acc. $\uparrow$
Qwen-2.5-7B-Instruct 18.1 31.6 34.4 62.1 62.0
gemini-2.0-flash 2.6 45.2 46.6 72.1 74.6
gemini-2.5-flash-preview-04-17 15.4 47.8 47.6 71.0 73.8
claude-3.5-sonnet-20241022 0.5 46.1 47.5 71.8 71.8
claude-3.7-sonnet-20250219 2.9 50.1 51.1 71.8 74.6
4.3 Ablation Study
Impact of LLMs on Code Generation. We evaluate the impact of using different LLMs on code generation quality in RefProg. We use Le3DE2E’s tracks for all experiments in this ablation. Somewhat unsurprisingly, we find that Claude 3.7 Sonnet performs the best (cf. Table 3), achieving a HOTA-Temporal of 50.1. We posit that this is because it has been explicitly tuned for code generation and instruction following. Gemini 2.0 Flash performs considerably worse despite also being tuned for coding. Interestingly, we find that Gemini and Qwen 2.5 Instruct struggle to generate valid programs, and instead try to generate full programs.
4.4 CVPR 2025 Scenario Mining Challenge
We are hosting a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining. To incentivize participation, top teams can win cash prizes. Although our competition will continue until June 8th 2025, we have already received submissions from three teams (some submissions are private).
4.5 Limitations and Future Work
Noisy Ground Truth Perception Labels. The quality of our scenario mining dataset is limited by the quality of the ground truth perception labels in AV2. For example, jittery tracks can lead to poor motion classification over short horizons. We address this issue by significantly post-processing and manually verifying all generated scenarios to mitigate label noise.
Limited Diversity. Although AV2 includes many interesting scenarios (cf. Fig. 4), it is still relatively small compared to industry-scale datasets. Moreover, the AV2 sensor dataset was primarily created to train and benchmark 3D perception models, and is not intended for scenario mining. As a result, it does not include many useful annotations like traffic light state, weather information, or lighting annotations. Although our dataset is of similar size to prior work like nuPrompt, future datasets should be explicitly curated to address scenario mining.
Runtime Complexity. RefProg requires running offline perception algorithms and executing LLM-based programs, both of which are prohibitively slow. Although these components are trivially parallelizable, we posit that our approach is not sufficiently efficient for real-world applications. Future work should also consider the runtime and scalability of their method beyond $1000$ logs.
Figure 4: Examples of Multi-Agent Interactions. We visualize representative examples from RefAV to highlight the diversity of our dataset. In (a), we capture the interactions between vulnerable road users and vehicles at a crowded intersection. Scenario (b) presents an atypical instance of a common multi-agent interaction (e.g. pedestrian walking a dog). In (c), we show a complex ego-vehicle trajectory that involves multiple moving vehicles. Scenario (d) illustrates an example of a rare multi-object interaction. In (e), we highlight a scenario that might requires evasive maneuvers from the ego-vehicle (e.g. the occluded pedestrian might cross the path of the ego-vehicle). Finally, subfigure (f) visualizes an example scene graph derived from AV2 annotations, where edges represent links between the referred object and related objects in the environment. Note that we show referred objects in green, related objects in blue, and all other objects in red.
5 Conclusion
In this paper, we introduce RefAV, a large-scale benchmark designed to evaluate scenario mining. Unlike prior language-based 3D scene understanding tasks, we find that scenario mining poses unique challenges in identifying complex multi-agent interactions. Notably prior referential tracking baselines struggle on this challenging benchmark, demonstrating the limitations of existing methods. Future work should develop models capable of reasoning over complex, multi-modal temporal data.
6 Acknowledgments
This work was supported in part by funding the NSF GRFP (Grant No. DGE2140739).
[…]
Appendix A Implementation Details
We present additional implementation details to reproduce our baseline experiments below. Our code is available on GitHub. We primarily consider two-stage baselines that first track all objects, then classify them into referred objects, related objects, or other objects. We repurpose 3D trackers from winning teams in the Argoverse 2 End-to-End Forecasting Challenge, and include a link to the base tracker outputs on Google Drive. All models are trained to output tracks for 26 annotated categories within Argoverse 2. We include an example prediction below:
1 {
2 "timestamp_ns": int,
3 "track_id": int,
4 "confidence": float,
5 "class_name": str,
6 "translation_m": np.array,
7 "size": np.array,
8 "yaw": np.array,
9 }
We append a track for the ego vehicle with a consistent track_id, confidence of 1.00, class name of EGO_VEHICLE, yaw of 0 radians, size (<length, width height> in meters) of <4.877, 2, 1.473>, and translation (<x, y, z> in meters) of <1.422, 0, 0.25>. This translation represents the offset from the centroid of the ego-vehicle to the ego-vehicle reference coordinate located near the center of its rear axle.
All Tracks Oracle. Our naive baseline labels all ground truth tracks as referred objects. This means that all logs and timestamps are labeled as “positive”, giving a balanced accuracy of 50%.
Filtering by Referred Class (RefBlind). RefBlind filters all tracks except those corresponding to the referred class. To get the referred class, we prompt Claude 3.7 Sonnet as follows:
Categories: {AV2 Categories} Please select the one category that most corresponds to the object of focus in the description:{Natural Language Description}. As an example, for the description "vehicle turning at intersection with nearby pedestrians" your output would be VEHICLE. For the description "ego vehicle near construction barrel" your output would be EGO_VEHICLE. Your only output should be the category name, in all capital letters and including underscores if necessary.
We define two super-categories for VEHICLE and ANY_OBJECT. The VEHICLE category is manually defined to contain the classes ARTICULATED_BUS, BOX_TRUCK, BUS, EGO_VEHICLE, LARGE_VEHICLE, MOTORCYCLE, RAILED_VEHICLE, REGULAR_VEHICLE, SCHOOL BUS, TRUCK, and TRUCK_CAB.
Referential Track Classification (RefClassify). RefClassify filters tracks by computing prompt-track CLIP similarity over time. At each timestep, we project the eight vertices of each object’s bounding box onto the image plane of all seven cameras. We then cull any objects that do not partially fall within the frustum of each camera. For each projected bounding box, if three or more vertices fall within the image plane, we take a crop using the minimum and maximum uv image coordinates. We pad this crop by 30 pixels on each side and feed the resulting image into a CLIP ViT-L/14 model to obtain CLIP image embeddings. We also use CLIP ViT-L/14 to compute text embeddings for all prompts in the dataset. Next, we calculate cosine similarity scores between each track’s image crops and the text prompt. Since individual similarity scores tend to be noisy, we follow ReferGPT [chamiti2025refergpt] in applying a dynamic threshold and majority voting strategy to identify the referred object tracks. Since the dataset includes both positive and negative examples for each prompt, we group similarity scores by prompt and threshold the top 10% of scores. If more than half of the scores of a track fall within the top 10% threshold, we label the entire track as a referred object. To reduce the number of length one tracks (i.e., tracks spanning only one timestamp), we apply a similarity score modifier: $s_{\text{modified}}=s_{\text{cosine}}-\frac{0.05}{len(\text{track})}$.
Referential Track Inflation (RefInflate). RefInflate leverages off-the-shelf foundational models to enable zero-shot 3D referential tracking. We implement a GroundingSAM-like pipeline [liu2024grounding, kirillov2023segment], where GroundingDino generates bounding boxes based on a given prompt, and SAM2 segments objects within those bounding boxes. We extend this process to predict 3D bounding boxes using multi-view video and LiDAR data [khurana2024shelf]. For each camera, we prompt GroundingDino with the input image and referential prompt, and extract bounding boxes that we then use to prompt SAM2 for object segmentation. We perform all operations at the native image resolution.
[…]
To associate bounding boxes across time, we use a greedy tracker that applies Hungarian Matching to match boxes from timestep $t$ to tracks from $t-1$ using world-coordinate distance. We set the maximum matching distance to $5$ meters.
Referential Tracking by Program Synthesis (RefProg). RefProg filters off-the-shelf tracks using Python programs synthesized by a large language model (LLM), without explicitly relying on any information from LiDAR or cameras. To reduce the time required to run these synthesized programs, we first filter the input tracks based on confidence. We compute each track’s confidence score by summing its confidence values across all timestamps. For the classes REGULAR_VEHICLE, PEDESTRIAN, BOLLARD, CONSTRUCTION_CONE, and CONSTRUCTION_BARREL, we retain the top 200 tracks per class. For all other classes, we retain the top 100 tracks per category. We prompt the LLM with the following prompt:
Please use the following functions to find instances of a referred object in an autonomous driving dataset. Be precise to the description, try to avoid returning false positives. API Listing: {API Listing} Categories: {AV2 Categories} Define a single scenario for the description:{Natural Language Description} Here is a list of examples: {Prediction Examples}. Only output code and comments as part of a Python block. Feel free to use a liberal amount of comments. Do not define any additional functions, or filepaths. Do not include imports. Assume the log_dir and output_dir variables are given. Wrap all code in one python block and do not provide alternatives. Output code even if the given functions are not expressive enough to find the scenario.
We include the API listing in Appendix F. We found this prompt consistently produces reasonable code across all tested LLMs. After running the synthesized programs, we post-process the output tracks in accordance with dataset practices. We remove all length one tracks as noise and discard object relationships where the spatial distance exceeds 50 meters.
For the remaining tracks, we apply timestamp dilation to reduce flickering of the referred object labels. We symmetrically dilate time segments to a minimum of 1.5 seconds. For example, if the base tracker tracks an object with ID 1 from 1.0 to 15.0 seconds at 2 Hz, and the synthesized code marks it as referred at timestamps 4.5 and 5.0, we update the final output to label the track as referred from 4.0 to 5.5 seconds. Finally, we downsample all output tracks to 2 Hz for evaluation.
Appendix B Scenario Mining Annotation Tool
In order to manually annotate interesting scenarios, we build a custom web app with Claude 3.7 (cf. Fig 5). To construct a scenario, the user selects an object by clicking on a point within the image frame, writes a natural language description, and selects the start and end frames of the referred objects correspond to the prompt. We project a ray originating from the 2D point and find the ground truth 3D bounding box centroid closest to the ray. This tool allows us to quickly generate multi-object and multi-camera referential tracks.
Figure 5: Manual Annotation Tool. We create an annotation tool to assist with labeling manually defined scenarios. Our tool allows us to quickly annotate multi-object referential tracks in AV2.
Appendix C Referential Tracking Evaluation of Related and Other Objects
RefProg creates a scene graph for each natural language prompt that not only identifies referred objects, but also classifies tracks as related objects and other objects. For example, given the prompt vehicle near a pedestrian, the vehicle would be the referred object, the pedestrian would be the related object, and everything else in the scene would be considered other objects. We benchmark the base tracker with HOTA and evaluate HOTA-Temporal of related objects and other objects in Table 4. Since most objects are not related to the prompt, HOTA and HOTA-Temporal Other are similar.
Table 4: Referential Tracking Accuracy for Related and Other Objects. We present RefProg’s referential tracking accuracy for related objects and other objects. We find that HOTA (for the base tracker) is similar to HOTA-Temporal (Other) because most objects are not relevant to the referential prompt.
            Method             HOTA $\uparrow$             HOTA-Temporal (Related) $\uparrow$             HOTA-Temporal (Other) $\uparrow$
            Referential Tracking by Program Synthesis
            Ground Truth             100.0             57.2             97.9
            Le3DE2E [chenle3de2e]             78.9             35.6             74.2
            ReVoxelDet [lee2024revoxeldet]             69.4             28.2             63.4
            TransFusion [bai2022transfusion]             69.5             28.6             63.8
            Valeo4Cast [xu2024valeo4cast]             75.1             30.8             69.2
Appendix D Impact of Sampling Rate on RefProg
We evaluate the impact of sampling rate on RefProg. First, we subsample three detectors (Le3DE2E [chenle3de2e], Valeo4Cast [xu2024valeo4cast], and BEVFusion [liu2022bevfusion]) at 2Hz and 10 Hz. Next, we associate all detections with AB3DMOT [weng20203d] to generate 2Hz and 10Hz tracks respectively. Lastly, we evaluate RefProg on all trackers at 2Hz. We find that standard tracking performance drops by 5% when subsampling due to the difficulty of associating sparse detections. We see a similar drop in HOTA-Track, Log Balanced Accuracy and Timestamp Balanced Accuracy. Interestingly, we do not see a similar drop in performance with HOTA-Temporal. This suggests that RefProg might be robust to temporally sparse inputs and ID switches.
Table 5: Impact of Sampling Rate. We evaluate RefProg’s performance with subsampled inputs to measure the relative impact on referential tracking accuracy. We find that standard tracking performance drops by 5% when subsampling due to the difficulty of associating sparse detections.
Method Hz HOTA $\uparrow$ HOTA-Temporal $\uparrow$ HOTA-Track $\uparrow$ Log Bal. Acc. $\uparrow$ Timestamp Bal. Acc. $\uparrow$
Ground Truth 10 100.0 64.7 68.7 81.1 80.7
Ground Truth 2 100.0 64.2 68.8 80.6 80.1
Le3DE2E Det. [chenle3de2e] + AB3DMOT [weng20203d] 10 78.0 48.3 49.2 73.1 73.6
Le3DE2E Det. [chenle3de2e] + AB3DMOT [weng20203d] 2 73.0 45.2 44.1 69.3 69.5
Valeo4Cast Det. [xu2024valeo4cast] + AB3DMOT [weng20203d] 10 74.8 42.8 45.5 71.6 71.8
Valeo4Cast Det. [liu2022bevfusion] + AB3DMOT [weng20203d] 2 69.9 45.2 44.1 69.3 69.5
BEVFusion Det. [liu2022bevfusion] + AB3DMOT [weng20203d] 10 76.0 46.4 47.8 73.7 72.3
BEVFusion Det. [liu2022bevfusion] + AB3DMOT [weng20203d] 2 70.9 46.4 45.2 72.1 72.9
Appendix E More Dataset Statistics
We present additional dataset statistics in Figure 6 and 7.
Expression Type # Expressions
Positive Prompts 7910
Negative Prompts 2090
Procedurally Defined 9880
Manually Defined 120
Includes Related Objects 5531
Only Ego is Referred 827
Only Ego is Related 1200
Figure 6: RefAV Expression Statistics. We compute the number of referring expressions categorized by expression type, count the number of the expressions that are positive or negative, and whether they are procedurally generated or manually constructed. We also highlight the number of positive expressions that include related object annotations and explicit references to the ego vehicle. The bar chart on the right shows the number of objects refered to by each expression. Most expressions in RefAV refer to zero (e.g. negative prompt) or one object. On average, each expression refers to 3.92 objects.
Figure 7: RefAV Object Distribution. RefAV includes referred objects in all directions up to 150m away from the ego vehicle. RefAV places a special focus on the ego vehicle and objects that interact with the ego vehicle. Therefore, referred objects are disproportionately located on the road in front of the ego vehicle. Both referred object heatmaps are in the ego vehicle coordinate frame.
Appendix F API Listing
We present our full Python API listing below.
1def has_objects_in_relative_direction(
2 track_candidates:dict,
3 related_candidates:dict,
4 log_dir:Path,
5 direction:Literal["forward", "backward", "left", "right"],
6 min_number:int=1,
7 max_number:int=np.inf,
8 within_distance:float=50,
9 lateral_thresh:float=np.inf) -> dict:
10 """
11 Identifies tracked objects with at least the minimum number of
12 related
13 candidates in the specified direction. If the minimum number is
14 met,
15 will create relationships equal to the max_number of closest
16 objects.
17
18 Args:
19 track_candidates: Tracks to analyze (scenario dictionary).
20 related_candidates: Candidates to check for in direction
21 (scenario dictionary).
22 log_dir: Path to scenario logs.
23 direction: Direction to analyze from the track’s point of view
24 (’forward’, ’backward’, ’left’, ’right’).
25 min_number: Minimum number of objects to identify in the
26 direction per timestamp. Defaults to 1.
27 max_number: Maximum number of objects to identify in the
28 direction per timestamp. Defaults to infinity.
29 within_distance: Maximum distance for considering an object in
30 the direction. Defaults to infinity.
31 lateral_thresh: Maximum lateral distance the related object
32 can be from the sides of the tracked object. Defaults to
33 infinity.
34
35 Returns:
36 dict:
37 A scenario dictionary where keys are track UUIDs and
38 values are dictionaries containing related candidate
39 UUIDs and lists of timestamps when the condition is
40 met for that relative direction.
41
42 Example:
43 vehicles_with_peds_in_front =
44 has_objects_in_relative_direction(vehicles, pedestrians,
45 log_dir, direction=’forward’, min_number=2)
46 """
47
48
49def get_objects_in_relative_direction(
50 track_candidates:dict,
51 related_candidates:dict,
52 log_dir:Path,
53 direction:Literal["forward", "backward", "left", "right"],
54 min_number:int=0,
55 max_number:int=np.inf,
56 within_distance:float=50,
57 lateral_thresh:float=np.inf)->dict:
58 """
59 Returns a scenario dictionary of the related candidates that are
60 in the relative direction of the track candidates.
61
62
63 Args:
64 track_candidates: Tracks (scenario dictionary).
65 related_candidates: Candidates to check for in direction
66 (scenario dictionary).
67 log_dir: Path to scenario logs.
68 direction: Direction to analyze from the track’s point of view
69 (’forward’, ’backward’, ’left’, ’right’).
70 min_number: Minimum number of objects to identify in the
71 direction per timestamp. Defaults to 0.
72 max_number: Maximum number of objects to identify in the
73 direction
74 per timestamp. Defaults to infinity.
75 within_distance: Maximum distance for considering an object in
76 the direction. Defaults to infinity.
77 lateral_thresh: Maximum lateral distance the related object
78 can be from the sides of the tracked object. Lateral distance
79 is distance is the distance from the sides of the object that
80 are parallel to the specified direction. Defaults to infinity.
81
82 Returns:
83 dict:
84 A scenario dictionary where keys are track UUIDs and
85 values are dictionaries containing related candidate UUIDs
86 and lists of timestamps when the condition is met for that
87 relative direction.
88
89 Example:
90 peds_in_front_of_vehicles =
91 get_objects_in_relative_direction(vehicles, pedestrians,
92 log_dir, direction=’forward’, min_number=2)
93 """
94
95
96def get_objects_of_category(log_dir, category)->dict:
97 """
98 Returns all objects from a given category from the log
99 annotations. This method accepts the super-categories "ANY" and "VEHICLE".
100
101 Args:
102 log_dir: Path to the directory containing scenario logs and
103 data.
104 category: the category of objects to return
105
106 Returns:
107 dict: A scenario dict that where keys are the unique id (uuid)
108 of the object and values are the list of timestamps the object
109 is in view of the ego-vehicle.
110
111 Example:
112 trucks = get_objects_of_category(log_dir, category=’TRUCK’)
113 """
114
115
116def is_category(track_candidates:dict, log_dir:Path, category:str):
117 """
118 Returns all objects from a given category from track_candidates
119 dict. This method accepts the super-categories "ANY" and "VEHICLE".
120
121 Args:
122 track_candidates: The scenario dict containing the objects to
123 filter down
124 log_dir: Path to the directory containing scenario logs and
125 data.
126 category: the category of objects to return
127
128 Returns:
129 dict: A scenario dict that where keys are the unique id of the
130 object of the given category and values are the list of
131 timestamps the object is in view of the ego-vehicle.
132
133 Example:
134 box_trucks = is_category(vehicles, log_dir,
135 category=’BOX_TRUCK’)
136 """
137
138
139def turning(
140 track_candidates: dict,
141 log_dir:Path,
142 direction:Literal["left", "right", None]=None)->dict:
143 """
144 Returns objects that are turning in the given direction.
145
146 Args:
147 track_candidates: The objects you want to filter from
148 (scenario dictionary).
149 log_dir: Path to scenario logs.
150 direction: The direction of the turn, from the track’s point
151 of view (’left’, ’right’, None).
152
153 Returns:
154 dict:
155 A filtered scenario dictionary where:
156 - Keys are track UUIDs that meet the turning criteria.
157 - Values are nested dictionaries containing timestamps.
158
159 Example:
160 turning_left = turning(vehicles, log_dir, direction=’left’)
161 """
162
163
164def changing_lanes(
165 track_candidates:dict,
166 log_dir:Path,
167 direction:Literal["left", "right", None]=None) -> dict:
168 """
169 Identifies lane change events for tracked objects in a scenario.
170
171 Args:
172 track_candidates: The tracks to analyze (scenario dictionary).
173 log_dir: Path to scenario logs.
174 direction: The direction of the lane change. None indicates
175 tracking either left or right lane changes
176 (’left’, ’right’, None).
177
178 Returns:
179 dict:
180 A filtered scenario dictionary where:
181 Keys are track UUIDs that meet the lane change criteria.
182 Values are nested dictionaries containing timestamps and
183 related data.
184
185 Example:
186 left_lane_changes = changing_lanes(vehicles, log_dir,
187 direction=’left’)
188 """
189
190
191def has_lateral_acceleration(
192 track_candidates:dict,
193 log_dir:Path,
194 min_accel=-np.inf,
195 max_accel=np.inf) -> dict:
196 """
197 Objects with a lateral acceleartion between the minimum and
198 maximum thresholds. Most objects with a high lateral acceleration
199 are turning. Postive values indicate accelaration to the left
200 while negative values indicate acceleration to the right.
201
202 Args:
203 track_candidates: The tracks to analyze (scenario dictionary).
204 log_dir: Path to scenario logs.
205 direction: The direction of the lane change. None indicates
206 tracking either left or right lane changes
207 (’left’, ’right’, None).
208
209 Returns:
210 dict:
211 A filtered scenario dictionary where:
212 Keys are track UUIDs that meet the lane change criteria.
213 Values are nested dictionaries containing timestamps and
214 related data.
215
216 Example:
217 jerking_left = has_lateral_acceleration(non_turning_vehicles,
218 log_dir, min_accel=2)
219 """
220
221
222def facing_toward(
223 track_candidates:dict,
224 related_candidates:dict,
225 log_dir:Path,
226 within_angle:float=22.5,
227 max_distance:float=50)->dict:
228 """
229 Identifies objects in track_candidates that are facing toward
230 objects in related candidates. The related candidate must lie
231 within a region lying within within_angle degrees on either
232 side the track-candidate’s forward axis.
233
234 Args:
235 track_candidates: The tracks that could be heading toward
236 another tracks
237 related_candidates: The objects to analyze to see if the
238 track_candidates are heading toward
239 log_dir: Path to the directory containing scenario logs
240 and data.
241 fov: The field of view of the track_candidates. The related
242 candidate must lie within a region lying within fov/2 degrees
243 on either side the track-candidate’s forward axis.
244 max_distance: The maximum distance a related_candidate can
245 be away to be considered by
246
247 Returns:
248 A filtered scenario dict that contains the subset of track
249 candidates heading toward at least one of the related
250 candidates.
251
252 Example:
253 pedestrian_facing_away = scenario_not(facing_toward)
254 (pedestrian, ego_vehicle, log_dir, within_angle=180)
255 """
256
257
258def heading_toward(
259 track_candidates:dict,
260 related_candidates:dict,
261 log_dir:Path,
262 angle_threshold:float=22.5,
263 minimum_speed:float=.5,
264 max_distance:float=np.inf)->dict:
265 """
266 Identifies objects in track_candidates that are heading toward
267 objects in related candidates. The track candidates acceleartion
268 vector must be within the given angle threshold of the relative
269 position vector. The track candidates must have a component of
270 velocity toward the related candidate greater than the
271 minimum_accel.
272
273 Args:
274 track_candidates: The tracks that could be heading toward
275 another tracks
276 related_candidates: The objects to analyze to see if the
277 track_candidates are heading toward
278 log_dir: Path to the directory containing scenario logs
279 and data.
280 angle_threshold: The maximum angular difference between
281 the velocity vector and relative position vector between
282 the track candidate and related candidate.
283 min_vel: The minimum magnitude of the component of velocity
284 toward the related candidate
285 max_distance: Distance in meters the related candidates can
286 be away from the track candidate to be considered
287
288 Returns:
289 A filted scenario dict that contains the subset of track
290 candidates heading toward at least one of the related
291 candidates.
292
293
294 Example:
295 heading_toward_traffic_cone = heading_toward(vehicles,
296 traffic_cone, log_dir)
297 """
298
299
300def accelerating(
301 track_candidates:dict,
302 log_dir:Path,
303 min_accel:float=.65,
304 max_accel:float=np.inf)->dict:
305 """
306 Identifies objects in track_candidates that have a forward
307 acceleration above a threshold. Values under -1 reliably
308 indicates braking. Values over 1.0 reliably indiciates
309 accelerating.
310
311 Args:
312 track_candidates: The tracks to analyze for acceleration
313 (scenario dictionary)
314 log_dir: Path to the directory containing scenario logs
315 and data.
316 min_accel: The lower bound of acceleration considered
317 max_accel: The upper bound of acceleration considered
318
319 Returns:
320 A filtered scenario dictionary containing the objects
321 with an acceleration between the lower and upper bounds.
322
323 Example:
324 accelerating_motorcycles = accelerating(motorcycles, log_dir)
325 """
326
327
328def has_velocity(
329 track_candidates:dict,
330 log_dir:Path,
331 min_velocity:float=.5,
332 max_velocity:float=np.inf)->dict:
333 """
334 Identifies objects with a velocity between the given maximum
335 and minimum velocities in m/s. Stationary objects may have a
336 velocity up to 0.5 m/s due to annotation jitter.
337
338 Args:
339 track_candidates: Tracks to analyze (scenario dictionary).
340 log_dir: Path to scenario logs.
341 min_velocity: Minimum velocity (m/s). Defaults to 0.5.
342 max_velocity: Maximum velocity (m/s)
343
344 Returns:
345 Filtered scenario dictionary of objects meeting the
346 velocity criteria.
347
348 Example:
349 fast_vehicles = has_min_velocity(vehicles, log_dir,
350 min_velocity=5)
351 """
352
353
354def at_pedestrian_crossing(
355 track_candidates:dict,
356 log_dir:Path,
357 within_distance:float=1)->dict:
358 """
359 Identifies objects that within a certain distance from a
360 pedestrian crossing.
361 A distance of zero indicates that the object is within the
362 boundaries of the pedestrian crossing.
363
364 Args:
365 track_candidates: Tracks to analyze (scenario dictionary).
366 log_dir: Path to scenario logs.
367 within_distance: Distance in meters the track candidate must
368 be from the pedestrian crossing. A distance of zero means
369 that the object must be within the boundaries of the
370 pedestrian crossing.
371
372 Returns:
373 Filtered scenario dictionary where keys are track UUIDs
374 and values are lists of timestamps.
375
376 Example:
377 vehicles_at_ped_crossing = at_pedestrian_crossing(vehicles,
378 log_dir)
379 """
380
381
382def on_lane_type(
383 track_uuid:dict,
384 log_dir,
385 lane_type:Literal["BUS", "VEHICLE", "BIKE"])->dict:
386 """
387 Identifies objects on a specific lane type.
388
389 Args:
390 track_candidates: Tracks to analyze (scenario dictionary).
391 log_dir: Path to scenario logs.
392 lane_type: Type of lane to check (’BUS’, ’VEHICLE’, or ’BIKE’).
393
394 Returns:
395 Filtered scenario dictionary where keys are track UUIDs
396 and values are lists of timestamps.
397
398 Example:
399 vehicles_on_bus_lane = on_lane_type(vehicles, log_dir,
400 lane_type="BUS")
401 """
402
403
404def near_intersection(
405 track_uuid:dict,
406 log_dir:Path,
407 threshold:float=5)->dict:
408 """
409 Identifies objects within a specified threshold of an intersection
410 in meters.
411
412 Args:
413 track_candidates: Tracks to analyze (scenario dictionary).
414 log_dir: Path to scenario logs.
415 threshold: Distance threshold (in meters) to define "near"
416 an intersection.
417
418 Returns:
419 Filtered scenario dictionary where keys are track UUIDs
420 and values are lists of timestamps.
421
422 Example:
423 bicycles_near_intersection = near_intersection(bicycles,
424 log_dir, threshold=10.0)
425 """
426
427
428def on_intersection(track_candidates:dict, log_dir:Path):
429 """
430 Identifies objects located on top of an road intersection.
431
432 Args:
433 track_candidates: Tracks to analyze (scenario dictionary).
434 log_dir: Path to scenario logs.
435
436 Returns:
437 Filtered scenario dictionary where keys are track UUIDs
438 and values are lists of timestamps.
439
440 Example:
441 strollers_on_intersection = on_intersection(strollers,
442 log_dir)
443 """
444
445
446def being_crossed_by(
447 track_candidates:dict,
448 related_candidates:dict,
449 log_dir:Path,
450 direction:Literal["forward", "backward", "left",
451 "right"]="forward",
452 in_direction:Literal[’clockwise’,’counterclockwise’,
453 ’either’]=’either’,
454 forward_thresh:float=10,
455 lateral_thresh:float=5)->dict:
456 """
457 Identifies objects that are being crossed by one of the
458 related candidate objects. A crossing is defined as the
459 related candidate’s centroid crossing the half-midplane
460 of a tracked candidate. The direction of the half-midplane
461 is specified with the direction.
462
463 Args:
464 track_candidates: Tracks to analyze .
465 related_candidates: Candidates (e.g., pedestrians
466 or vehicles) to check for crossings.
467 log_dir: Path to scenario logs.
468 direction: specifies the axis and direction the
469 half midplane extends from
470 in_direction: which direction the related candidate has
471 to cross the midplane for it to be considered a crossing
472 forward_thresh: how far the midplane extends from the edge of
473 the tracked object
474 lateral_thresh: the two planes offset from the midplane. If an
475 related candidate crosses the midplane, it will continue being
476 considered crossing until it goes past the lateral_thresh.
477
478 Returns:
479 A filtered scenario dictionary containing all of the track
480 candidates that were crossed by the related candidates given
481 the specified constraints.
482
483 Example:
484 overtaking_on_left = being_crossed_by(moving_cars,
485 moving_cars, log_dir, direction="left", in_direction=
486 "clockwise", forward_thresh=4)
487 vehicles_crossed_by_peds = being_crossed_by(vehicles,
488 pedestrians, log_dir)
489 """
490
491
492def near_objects(
493 track_uuid:dict,
494 candidate_uuids:dict,
495 log_dir:Path,
496 distance_thresh:float=10,
497 min_objects:int=1,
498 include_self:bool=False)->dict:
499 """
500 Identifies timestamps when a tracked object is near a specified
501 set of related objects.
502
503 Args:
504 track_candidates: Tracks to analyze (scenario dictionary).
505 related_candidates: Candidates to check for proximity
506 (scenario dictionary).
507 log_dir: Path to scenario logs.
508 distance_thresh: Maximum distance in meters a related
509 candidate can be away to be considered "near".
510 min_objects: Minimum number of related objects required to be
511 near the tracked object.
512
513 Returns:
514 dict:
515 A scenario dictionary where:
516 Keys are timestamps when the tracked object is near the
517 required number of related objects.
518 Values are lists of related candidate UUIDs present at
519 those timestamps.
520
521 Example:
522 vehicles_near_ped_group = near_objects(vehicles, pedestrians,
523 log_dir, min_objects=3)
524 """
525
526
527def following(
528 track_uuid:dict,
529 candidate_uuids:dict,
530 log_dir:Path) -> dict:
531 """
532 Returns timestamps when the tracked object is following a lead
533 object
534 Following is defined simultaneously moving in the same direction
535 and lane.
536 """
537
538
539def heading_in_relative_direction_to(track_candidates,
540related_candidates, log_dir, direction:Literal[’same’,
541’opposite’, ’perpendicular’]):
542 """
543 Returns the subset of track candidates that are traveling
544 in the given direction compared to the related candidates.
545
546 Arguements:
547 track_candidates: The set of objects that could be traveling
548 in the given direction
549 related_candidates: The set of objects that the direction is
550 relative to
551 log_dir: The path to the log data
552 direction: The direction that the positive tracks are
553 traveling in relative to the related candidates
554 "opposite" indicates the track candidates are traveling in
555 a direction
556 135-180 degrees from the direction the related candidates
557 are heading toward.
558 "same" indicates the track candidates that are traveling
559 in a direction
560 0-45 degrees from the direction the related candiates
561 are heading toward.
562 "same" indicates the track candidates that are traveling
563 in a direction
564 45-135 degrees from the direction the related candiates
565 are heading toward.
566
567 Returns:
568 the subset of track candidates that are traveling in the
569 given direction compared to the related candidates.
570
571 Example:
572 oncoming_traffic = heading_in_relative_direction_to(vehicles,
573 ego_vehicle, log_dir, direction=’opposite’)
574 """
575
576
577def stationary(track_candidates:dict, log_dir:Path):
578 """
579 Returns objects that moved less than 2m over their length of
580 observation in the scneario. This object is only intended to
581 separate parked from active vehicles. Use has_velocity() with
582 thresholding if you want to indicate vehicles that are
583 temporarily stopped.
584
585 Args:
586 track_candidates: Tracks to analyze (scenario dictionary).
587 log_dir: Path to scenario logs.
588
589 Returns:
590 dict:
591 A filtered scenario dictionary where keys are track UUIDs
592 and values are lists of timestamps when the object is
593 stationary.
594
595 Example:
596 parked_vehicles = stationary(vehicles, log_dir)
597 """
598
599
600def at_stop_sign(track_candidates:dict, log_dir:Path, forward_thresh:float=10):
601 """
602 Identifies timestamps when a tracked object is in a lane
603 corresponding to a stop sign. The tracked object must be
604 within 15m of the stop sign. This may highlight vehicles
605 using street parking near a stopped sign.
606
607 Args:
608 track_candidates: Tracks to analyze (scenario dictionary).
609 log_dir: Path to scenario logs.
610 forward_thresh: Distance in meters the vehcile is from
611 the stop sign in the stop sign’s front direction
612
613 Returns:
614 dict:
615 A filtered scenario dictionary where keys are track UUIDs
616 and values are lists of timestamps when the object is at
617 a stop sign.
618
619 Example:
620 vehicles_at_stop_sign = at_stop_sign(vehicles, log_dir)
621 """
622
623
624def in_drivable_area(track_candidates:dict, log_dir:Path)->dict:
625 """
626 Identifies objects within track_candidates that are within a
627 drivable area.
628
629 Args:
630 track_candidates: Tracks to analyze (scenario dictionary).
631 log_dir: Path to scenario logs.
632
633 Returns:
634 dict:
635 A filtered scenario dictionary where keys are track UUIDs
636 and values are lists of timestamps when the object is in
637 a drivable area.
638
639 Example:
640 buses_in_drivable_area = in_drivable_area(buses, log_dir)
641 """
642
643
644def on_road(
645 track_candidates:dict,
646 log_dir:Path)->dict:
647 """
648 Identifies objects that are on a road or bike lane.
649 This function should be used in place of in_driveable_area() when
650 referencing objects that are on a road. The road does not include
651 parking lots or other driveable areas connecting the road to
652 parking lots.
653
654 Args:
655 track_candidates: Tracks to filter (scenario dictionary).
656 log_dir: Path to scenario logs.
657
658 Returns:
659 The subset of the track candidates that are currently on a road.
660
661 Example:
662 animals_on_road = on_road(animals, log_dir)
663 """
664
665
666def in_same_lane(
667 track_candidates:dict,
668 related_candidates:dict,
669 log_dir:Path) -> dict:
670 """
671 "
672 Identifies tracks that are in the same road lane as a related
673 candidate.
674
675 Args:
676 track_candidates: Tracks to filter (scenario dictionary)
677 related_candidates: Potential objects that could be in
678 the same lane as the track (scenario dictionary)
679 log_dir: Path to scenario logs.
680
681 Returns:
682 dict:
683 A filtered scenario dictionary where keys are track
684 UUIDs and values are lists of timestamps when the
685 object is on a road lane.
686
687 Example:
688 bicycle_in_same_lane_as_vehicle = in_same_lane(bicycle,
689 regular_vehicle, log_dir)
690 """
691
692
693def on_relative_side_of_road(
694 track_candidates:dict,
695 related_candidates:dict,
696 log_dir:Path,
697 side=Literal[’same’, ’opposite’]) -> dict:
698 """
699 "
700 Identifies tracks that are in the same road lane as a related
701 candidate.
702
703 Args:
704 track_candidates: Tracks to filter (scenario dictionary)
705 related_candidates: Potential objects that could be in
706 the same lane as the track (scenario dictionary)
707 log_dir: Path to scenario logs.
708
709 Returns:
710 dict:
711 A filtered scenario dictionary where keys are track UUIDs
712 and values are lists of timestamps when the object is on
713 a road lane.
714
715 Example:
716 bicycle_in_same_lane_as_vehicle = in_same_lane(bicycle,
717 regular_vehicle, log_dir)
718 """
719
720
721def scenario_and(scenario_dicts:list[dict])->dict:
722 """
723 Returns a composed scenario where the track objects are the
724 intersection of all of the track objects with the same uuid
725 and timestamps.
726
727 Args:
728 scenario_dicts: the scenarios to combine
729
730 Returns:
731 dict:
732 a filtered scenario dictionary that contains tracked
733 objects found in all given scenario dictionaries
734
735 Example:
736 jaywalking_peds = scenario_and([peds_on_road,
737 peds_not_on_pedestrian_crossing])
738 """
739
740
741def scenario_or(scenario_dicts:list[dict]):
742 """
743 Returns a composed scenario where that tracks all objects and
744 relationships in all of the input scenario dicts.
745
746 Args:
747 scenario_dicts: the scenarios to combine
748
749 Returns:
750 dict:
751 an expanded scenario dictionary that contains every
752 tracked object in the given scenario dictionaries
753
754 Example:
755 be_cautious_around = scenario_or([animal_on_road,
756 stroller_on_road])
757 """
758
759
760def reverse_relationship(func):
761 """
762 Wraps relational functions to switch the top level tracked objects
763 and relationships formed by the function.
764
765 Args:
766 relational_func: Any function that takes track_candidates and
767 related_candidates as its first and second arguements
768
769 Returns:
770 dict:
771 scenario dict with swapped top-level tracks and related
772 candidates
773
774 Example:
775 group_of_peds_near_vehicle =
776 reverse_relationship(near_objects)(vehicles, peds,
777 log_dir, min_objects=3)
778 """
779
780
781def scenario_not(func):
782 """
783 Wraps composable functions to return the difference of the input
784 track dict and output scenario dict. Using scenario_not with a
785 composable relational function will not return any relationships.
786
787 Args:
788 composable_func: Any function that takes track_candidates as
789 its first input
790
791 Returns:
792
793 Example:
794 active_vehicles = scenario_not(stationary)(vehicles, log_dir)
795 """
796
797
798def output_scenario(
799 scenario:dict,
800 description:str,
801 log_dir:Path,
802 output_dir:Path,
803 visualize:bool=False,
804 **visualization_kwargs):
805 """
806 Outputs a file containing the predictions in an evaluation-ready
807 format. Do not provide any visualization kwargs.
808 """

### v2

@@
 Notably, we find that naively repurposing off-the-shelf VLMs yields poor performance, suggesting that scenario mining presents unique challenges.
+Lastly, we discuss our recent CVPR 2025 competition and share insights from the community.
 Our code and dataset are available on Github and Argoverse.
@@
 Our extensive experiments highlight the limitations of current methods, and demonstrate the effectiveness of our proposed program synthesis-based approach.
-Lastly, we host a challenge at CVPR 2025 in conjunction with the Workshop on Autonomous Driving to encourage broad community involvement in addressing this challenging problem.
+Lastly, we highlight the results of our recent CVPR 2025 challenge hosted in conjunction with the Workshop on Autonomous Driving.
 2 Related Works
@@
 Figure 2 left).
-We include a full API listing in Appendix F.
+We include a full API listing in Appendix G.
 To generate a new scenario definition, we provide an LLM (e.g.
@@
 Valeo4Cast [xu2024valeo4cast] 69.2 41.7 44.1 62.7 69.1
+Challenge Submissions
+Zeekr UMCV 74.4 53.4 51.0 66.3 76.6
+Mi3 UCM 74.4 52.4 51.5 65.8 77.5
+ZXH 74.4 52.1 50.2 66.5 76.1
 4.2 Empirical Analysis of Results
@@
 4.4 CVPR 2025 Scenario Mining Challenge
-We are hosting a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining.
-To incentivize participation, top teams can win cash prizes.
-Although our competition will continue until June 8th 2025, we have already received submissions from three teams (some submissions are private).
+We hosted a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining.
+Our competition recieved submissions from eight teams (some submissions are private) at the close of our competition on June 8th, 2025 AOE.
+Notably, four teams beat our best baseline.
+We present the current top three teams at the bottom of Table 2.
+We summarize the contributions of the top three teams in Appendix F and include a link to full technical reports and code here.
 4.5 Limitations and Future Work
@@
 Scenario (d) illustrates an example of a rare multi-object interaction.
-In (e), we highlight a scenario that might requires evasive maneuvers from the ego-vehicle (e.g.
+In (e), we highlight a scenario that might require evasive maneuvers from the ego-vehicle (e.g.
 the occluded pedestrian might cross the path of the ego-vehicle).
-Finally, subfigure (f) visualizes an example scene graph derived from AV2 annotations, where edges represent links between the referred object and related objects in the environment.
+Finally, subfigure (f) visualizes a scenario with a multiple-step relationship (e.g.
+the official signaler is standing inside of a construction zone).
 Note that we show referred objects in green, related objects in blue, and all other objects in red.
@@
 6 Acknowledgments
-This work was supported in part by funding the NSF GRFP (Grant No. DGE2140739).
+This work was supported in part by funding from Bosch Research and the NSF GRFP (Grant No. DGE2140739). We would also like to thank Ishan Khatri for his feedback on early drafts
 […]
@@
 Output code even if the given functions are not expressive enough to find the scenario.
-We include the API listing in Appendix F.
+We include the API listing in Appendix G.
 We found this prompt consistently produces reasonable code across all tested LLMs.
@@
 Both referred object heatmaps are in the ego vehicle coordinate frame.
-Appendix F API Listing
+Appendix F Summary of CVPR 2025 Competition Top Performers
+We summarize the contributions of top teams below. We present full technical reports and code here.
+Zeekr UMCV extends RefProg with global context-aware generation to processes all queries collectively for improved reasoning and consistency.
+They also propose a multi-agent refinement system where a secondary refiner agent iteratively debugs and enhances the generated code.
+Mi3 UCM proposes two key improvements to RefProg: (1) Fault-Tolerant Iterative Code Generation (FT-ICG), which refines faulty code by re-prompting the LLM with error feedback until successful execution, and (2) Enhanced Prompting for Spatial Relational Functions (EP-SRF), which clarifies function semantics to prevent misinterpretations of spatial relationships between objects.
+ZXH refines RefProg by prompting LLMs with a batch of queries and in-context examples instead of prompting for each query one at a time.
+Importantly, ZXH’s approach relies on in-context learning and batch prompting to improve consistency and efficiency.
+Appendix G API Listing
 We present our full Python API listing below.